OpenStudy (anonymous):

Angelo took a group of people to the movies and spent $80 on tickets. Adult tickets cost $10 and child tickets cost $6. The equation to represent his cost is 10A + 6C = $80. If there were 5 children in the group, how many adult tickets did he purchase? Show your work and use a complete sentence to explain your solution in the context of the original problem , please help me step by step

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

10A + 6C = $80

ganeshie8 (ganeshie8):

you're given that there were 5 children in the group

ganeshie8 (ganeshie8):

so simply plugin C = 5 and solve for A

ganeshie8 (ganeshie8):

10A + 6*5 = 80 10A + 30 = 80

ganeshie8 (ganeshie8):

solve A

OpenStudy (anonymous):

how do I solve A?

OpenStudy (debbieg):

First you need to isolate the term that contains the A, so "move the 30" to the other side by subtracting it from both sides. What does that give you?

OpenStudy (anonymous):

what am I subtracting 30 too?:O

OpenStudy (debbieg):

It's an equation. You can apply the equality properties to an equation - add/subtract the same number from each side. Multiply/divide on each side by the same non-zero number. Use these to solve equations (you should already have covered basic equation solving techniques, given this problem). E.g.: if: 2x+10=14 subtract 10 from each side 2x+10-10=14-10 2x=4 Now you try it with 10A + 30 = 80 Subtract 30 from each side. What do you get? (do JUST that step)

OpenStudy (anonymous):

a=50?

OpenStudy (anonymous):

@DebbieG

OpenStudy (debbieg):

Not A=50..... 10A+30=80 10A+30-30=80-30 10A=50 Where did your 10 go?

OpenStudy (anonymous):

oh I understand !

OpenStudy (anonymous):

is that all I do?

OpenStudy (debbieg):

No, that was the first step. So now you have: 10A=50 Then what does A=??

OpenStudy (anonymous):

a=5

OpenStudy (debbieg):

very good. :)

OpenStudy (anonymous):

thank you so muchhh(: